# An Efficient Sign Language Translation Using Spatial Configuration and Motion Dynamics with LLMs

### v1

@@ 5 Conclusion
In this paper, we introduce SpaMo, a novel gloss-free SLT framework based on LLMs. Apart from the previous methods that rely on domain-specific fine-tuning of their visual encoders, SpaMo focuses on capturing the spatial configurations and motion dynamics, eliminating the need for resource-intensive fine-tuning. We also propose VT-Align, a training strategy that effectively aligns and narrows the modality gap between the sign videos and target translations, enabling the transformation of the sign videos into inputs interpretable by the LLM. Our approach achieves state-of-the-art results on two widely-used datasets. Furthermore, we provide the first comprehensive analysis of how the LLM interpret the sign videos within its embedding space and translate them into corresponding text.
@@
Limited Use of LLMs.
Currently, SpaMo has been tested on a limited range of LLMs. As shown in Table 4.3, our method scales effectively as the size of the LLMs increases. Therefore, there is a significant opportunity to expand this testing to include various other models, such as Llama Touvron et al. (2023), Vicuna Chiang et al. (2023), and Alpaca Taori et al. (2023). The primary constraint has been the limited availability of GPU resources. Expanding the testing to more LLMs could provide deeper insights and potentially enhance SpaMo’s performance across different architectures. Additionally, techniques such as 4-bit quantization can be employed to optimize these models, reducing the computational resources required and enabling more extensive testing. Future work will focus on broadening the range of tested models and exploring these optimization techniques to further improve the scalability and efficiency of SpaMo across diverse LLMs.
Limited Use of Sign Language Datasets.
In recent studies Uthus et al. (2024); Rust et al. (2024), scaling datasets has consistently led to performance improvements, as seen with larger sign language datasets, such as Youtube-ASL Uthus et al. (2024). While the dataset scaling could also enhance our method, in this work, we focus on a constrained setting. Specifically, we use a limited sign language dataset to evaluate and compare results, demonstrating the effectiveness of our method in resource-limited scenarios. Future work will involve expanding the dataset size to explore the full potential of our method and to assess its scalability and performance across more extensive and diverse datasets.
LoRA Fine-Tuning for LLM.
In this paper, we highlight that domain-specific fine-tuning of visual encoders is not essential for our method. However, our method incorporates LoRA fine-tuning for the LLM. While this might appear to be a compromise, it significantly reduces the resource requirements compared to fine-tuning both the visual encoders and the LLM. Additionally, as we discussed in the previous section, this limitation can be addressed as more data becomes available, allowing for improved scalability and performance over time.
@@
Our work is focused on developing a practical framework for sign language translation with the goal of overcoming communication barriers faced by the Deaf and hard-of-hearing communities. Although our approach utilizes off-the-shelf visual encoders and LLMs, there is a possibility that the framework could produce unexpected or biased outputs due to the inherent limitations in the pre-trained models. However, we are optimistic that future advancements in LLMs will help mitigate these issues. We rely on open datasets such as PHOENIX14T Camgoz et al. (2018) and How2Sign Duarte et al. (2021), which, while containing potentially identifiable information, present minimal concerns regarding personal privacy. Additionally, our method has been validated only on American and German sign languages, limiting its applicability to other sign languages. We call for future research in sign language translation to expand the diversity of sign language datasets, such as YouTube-ASL Uthus et al. (2024), BOBSL Albanie et al. (2021), and CSL-Daily Zhou et al. (2021a), to enhance the framework’s applicability and inclusively across different sign languages.
@@
In this Supplementary Material, we begin by providing further implementation details in Section A. Section B presents additional experimental results. In Section C, we delve into a more detailed analysis, offering statistics and insights on the sign language datasets used in our study. Finally, in Section D, we conclude by demonstrating additional qualitative results for PHOENIX14T and How2Sign.
@@ Appendix A More Implementation Details
Prompt Template.
To focus the LLM on the SLT task, we employ a specific prompting strategy. Our prompt includes a clear instructive prompt: “Translate the given sentence into German.” Following this, we incorporate multilingual translations via a translation engine such as Google Translator [^5] https://cloud.google.com/translate, which are sampled from the training set. These translations are included to facilitate In-Context Learning (ICL) Brown et al. (2020). The prompt is structured as follows: “Translate the given sentence into German. [SRC] = [TRG].” Here, the source input (e.g., a sentence in French) serves as the foreign language example, and the corresponding response is the translation into the target language (e.g., German, as used in PHOENIX14T). An example of this prompt structure is provided in Table A. To ensure that the LLM does not directly access the target translations during training, we shuffle the translation samples so they do not match the target translation. At test time, we select a translation pair from the training set to use as a reference.
@@
PHOENIX14Camgoz et al. (2018) DGS 3K 7,096 / 519 / 642 116 O Weather Forecast
How2Sign Duarte et al. (2021) ASL 16K 31,128 / 1,741 / 2,322 173 X Instructional
Table 11: Statistics of two sign language datasets used in this work. DGS: German Sign Language; ASL: American Sign Language; Avg. No. Frame: average number of video frames.
Training.
For training, we use the AdamW optimizer Loshchilov and Hutter (2017), with $\beta_{1}=0.9$, $\beta_{2}=0.98$, and a weight decay of $0.01$. The learning rate schedule includes a cosine decay with a peak learning rate of 1e-4 and a linear warmup of than 10k steps, with a minimum learning rate of 5e-5. We train our model for 40 epochs, using a single NVIDIA A100 GPU, completing the entire process within 24 hours.
Evaluating Process with KDEs.
To evaluate the quality of the learned representations, we utilize Kernel Density Estimation (KDE) to estimate the probability density functions of the embeddings from GFSLT-VLP and ours. Due to different dimensionality from these methods (1,024 vs. 2,048), we run Principal Component Analysis (PCA) to reduce the number of dimensions while retaining the most significant variance components. This dimensionality reduction facilitated more efficient and stable KDE fitting. KDE can be expressed as:
@@
Appendix B More Experiments
Effect of Visual Encoders.
We assess the effect of various combination of visual encoders (SE & ME). Table A shows four different encoders: DINOv2 Oquab et al. (2023), CLIP Radford et al. (2021), V-JEPA Bardes et al. (2024), and VideoMAE Tong et al. (2022). The results demonstrate that the combination of CLIP and VideoMAE delivers the highest performance, suggesting potential for further improvement as visual encoders continue to advance.
Effect of LoRA.
@@ Appendix B More Experiments
Appendix C Statistics of Sign Language Datasets
Table 11 provides an overview of two sign language datasets: PHOENIX14T and How2Sign, which differ significantly in their characteristics and applications. PHOENIX14T focuses on German Sign Language (DGS) within the specific domain of weather forecasting, featuring a relatively small vocabulary of 3K words and a concise average video length of 116 frames. It includes 7,096 training samples, 519 validation samples, and 642 test samples, with gloss annotations available. This dataset is tailored for domain-specific tasks, offering clear and repetitive patterns ideal for translation and recognition within weather-related contexts.
In contrast, How2Sign, centered on American Sign Language (ASL) in the instructional domain, presents a much larger and more diverse dataset with a 16K word vocabulary and an average of 173 frames per video. It provides a substantial amount of data, with 31,128 training samples, 1,741 validation samples, and 2,322 test samples, though it lacks gloss annotations. The broader and more complex nature of How2Sign makes it suitable for general sign language processing tasks, especially those requiring an understanding of diverse and intricate sign sequences.
@@ Appendix C Statistics of Sign Language Datasets
Appendix D More Qualitative Results
We provide additional translation examples for PHOENIX14T and How2Sign in Tables 12 and 13, respectively. As shown in Table 12, our method consistently delivers accurate translations, while GFSLT-VLP struggles to capture the correct semantic meaning.
For How2Sign, Table 13 presents translation results along with their corresponding visual tokens. Since How2Sign lacks gloss annotations, we include actual sign frames for qualitative comparison. Similar to the PHOENIX14T results, several visual tokens in How2Sign are closely aligned with the translations. Note that although OpenSLT Tarrés et al. (2023) is the only publicly available baseline [^6] https://github.com/imatge-upc/slt_how2sign_wicv2023, we were unable to reproduce their results due to a broken link to the fine-tuned I3D features at the time of drafting.

### v2

@@ 5 Conclusion
In this paper, we introduced SpaMo, a novel gloss-free SLT framework based on LLMs. Apart from the previous methods that rely on domain-specific fine-tuning of their visual encoders, SpaMo focuses on capturing the spatial configurations and motion dynamics, eliminating the need for resource-intensive fine-tuning. We also proposed VT-Align, a training strategy that effectively aligns and narrows the modality gap between the sign videos and target texts, enabling the transformation of the sign videos into inputs interpretable by the LLM. Our approach achieved state-of-the-art results on three popular datasets. Furthermore, we provided the first comprehensive analysis of how the LLM interprets the sign videos within its embedding space and translates them into corresponding text.
@@
Recently, scaling datasets Uthus et al. (2024); Rust et al. (2024) has consistently led to performance improvements, as seen with larger sign language datasets, such as Youtube-ASL Uthus et al. (2024) and BOBSL Albanie et al. (2021). While dataset scaling could also enhance our method, in this work, we focus on a constrained setting. Specifically, we use limited sign language datasets to evaluate and compare results, demonstrating the effectiveness of our method in resource-limited scenarios. Future work will involve expanding the dataset size to explore the full potential of our method and to assess its scalability and performance across more extensive and diverse datasets.
In this paper, we highlight that domain-specific fine-tuning of visual encoders is not essential for our method. However, this implies that our method relies on visual encoders pre-trained on general tasks such as action recognition and image captioning. To bridge this gap, we introduce a pre-alignment process and apply LoRA fine-tuning to the LLM. While this might appear to be a compromise, it significantly reduces the resource requirements compared to fine-tuning both the visual encoders and the LLM. Additionally, as we discussed in the previous paragraph, this limitation can be addressed as more data becomes available, allowing for improved scalability and performance over time.
@@
Our work focuses on developing a practical framework for SLT with the goal of overcoming communication barriers faced by the Deaf and hard-of-hearing communities. Although our approach utilizes off-the-shelf visual encoders and LLMs, there is a possibility that the framework could produce unexpected or biased outputs due to the inherent limitations in the pre-trained models. However, we are optimistic that future advancements in LLMs will help mitigate these issues. We rely on open datasets such as PHOENIX14T Camgoz et al. (2018), CSL-Daily Zhou et al. (2021a), and How2Sign Duarte et al. (2021), which, while containing potentially identifiable information, present minimal concerns regarding personal privacy. Additionally, our method has been validated only on German, Chinese, and American sign languages, limiting its applicability to other sign languages. We call for future research in SLT to expand on a broader range of sign languages, promoting greater equity for the Deaf community.
Acknowledgment
This work was supported by the Institute for Information and communications Technology Promotion (IITP) grant funded by the Korea government (MSIT) (No. 2022-0-00010, Development of Korean sign language translation service technology for the deaf in medical environment).
@@ Acknowledgment
In this Appendix, we first provide additional implementation details in Section A. Then, Section B provides more details about the sign language dataset used in this study, including its statistics. In Section C, we present further experimental results. Finally, in Section D, we discuss the feasibility of existing Vision-Language Models (VLMs) in the SLT domain.
@@ Appendix A More Implementation Details
A.2 Prompt Template.
To focus the LLM on the SLT task, we employ a specific prompting strategy. Our prompt includes a clear instructive prompt: “Translate the given sentence into German.” Following this, we incorporate multilingual translations via a translation engine such as Google Translate [^5] https://cloud.google.com/translate, which are sampled from the training set. These translations are included to facilitate In-Context Learning (ICL) Brown et al. (2020). The prompt is structured as follows: “Translate the given sentence into German. [SRC] = [TRG].” Here, the source input (e.g., a sentence in French) serves as the foreign language example, and the corresponding response is the translation into the target language (e.g., German, as used in PHOENIX14T). An example of this prompt structure is provided in Table A.2. To ensure that the LLM does not directly access the target translations during training, we shuffle the translation samples so that they do not match the target translation. At test time, we select a translation pair from the training set to use as a reference.
@@
PHOENIX14T Camgoz et al. (2018) DGS 3K 7,096 / 519 / 642 116 $\checkmark$ Weather Forecast
CSL-Daily Zhou et al. (2021a) CSL 2K 18,401 / 1,077 / 1,176 119 $\checkmark$ Daily-life
How2Sign Duarte et al. (2021) ASL 16K 31,128 / 1,741 / 2,322 173 Instructional
Table 11: Statistics of three sign language datasets used in this work. DGS: German Sign Language; CSL: Chinese Sign Language; ASL: American Sign Language; Avg. No. Frame: average number of video frames.
A.3 Training.
For training, we use the AdamW optimizer Loshchilov and Hutter (2017), with $\beta_{1}=0.9$, $\beta_{2}=0.98$, and a weight decay of $0.01$. The learning rate schedule includes a cosine decay with a peak learning rate of 1e-4 and a linear warmup of over 10K steps, with a minimum learning rate of 5e-5. We train our model for 40 epochs, using a single NVIDIA A100 GPU, completing the entire process within 24 hours.
A.4 Evaluating Process with KDEs.
To evaluate the quality of the learned representations, we utilize Kernel Density Estimation (KDE) to estimate the probability density functions of the embeddings from GFSLT-VLP and ours. Due to different dimensionality between these methods (1,024 vs. 2,048), we run Principal Component Analysis (PCA) to reduce the number of dimensions while retaining the most significant variance components. This dimensionality reduction facilitates more efficient and stable KDE fitting. KDE can be expressed as:
@@
A.5 Generating Visual Tokens
Inspired by the reverse engineering Ju et al. (2023), we first compute the Euclidean distance between the sign feature $Z_{sm}$ and the LLM’s embedding table $E_{llm}\in\mathbb{R}^{V\times d^{\prime}}$, where $V$ represents the vocabulary size. Each sign feature is then mapped to the word associated with the shortest distance in this space. This process can be expressed as $\text{dist}(Z_{sm},E_{llm})\leq\Delta$, where $\text{dist}(\cdot)$ denotes the Euclidean distance function, and $\Delta$ represents the shortest distance to $E_{llm}$ across all sign features.
Appendix B Statistics of Sign Language Datasets
Table 11 presents a comparative overview of three popular sign language datasets: PHOENIX14T, CSL-Daily, and How2Sign, each with distinct statistics and domain.
PHOENIX14T focuses on German Sign Language (DGS) within the specific domain of weather forecasting, featuring a relatively small vocabulary of 3K words and a concise average video length of 116 frames. It includes 7,096 training samples, 519 validation samples, and 642 test samples, with gloss annotations available. This dataset is tailored for domain-specific tasks, offering clear and repetitive patterns ideal for translation and recognition within weather-related contexts.
In comparison, CLS-Daily, a dataset for Chinese Sign Language (CSL), covers a broader range of topics than PHOENIX14T, spanning areas such as family life, medical care, school, banking, shopping, and social interactions. It features a vocabulary of 2K words and an average video length of 119 frames. The dataset includes 18,401 training samples, 1,077 validation samples, and 1,176 test samples, also with gloss annotations.
On the other hand, How2Sign focuses on American Sign Language (ASL) in the instructional domain. It offers a significantly larger and more diverse dataset, with a vocabulary of 16K words and an average video length of 173 frames. The dataset consists of 31,128 training samples, 1,741 validation samples, and 2,322 test samples, but lacks gloss annotations. The diversity and complexity of How2Sign make it particularly suitable for general sign language related tasks, especially those that involve understanding varied and intricate sign sequences.
Appendix C More Experiments
C.1 Effect of Visual Encoders.
We assess the effect of various combinations of visual encoders (SE & ME). Table A.2 shows four different encoders: DINOv2 Oquab et al. (2023), CLIP Radford et al. (2021), V-JEPA Bardes et al. (2024), and VideoMAE Tong et al. (2022). The results demonstrate that the combination of CLIP and VideoMAE delivers the highest performance, suggesting potential for further improvement as visual encoders continue to advance.
C.2 Effect of LoRA.
@@ Appendix C More Experiments
Methods Vis. Ft. #Trainable Params #Total Params B4
GFSLT-VLP Zhou et al. (2023) $\checkmark$ 215.6M 215.6M 21.44
Sign2GPT Wong et al. (2024) $\checkmark$ 16M 1.8B 22.52
Fla-LLM Chen et al. (2024) $\checkmark$ >705.6M* >705.6M* 23.09
SignLLM Gong et al. (2024) $\checkmark$ - >7B* 23.40
[…]
Table 12: Model parameter comparison. * denotes an estimated value due to the unavailability of public code. “Vis. Ft.” denotes to the visually fine-tuned on sign language datasets.
C.3 Parameter Comparisons
We present a comparison of various SLT methods, focusing on the presence of visual fine-tuning ("Vis. Ft."), the number of trainable and total parameters, and their performance as measured by the BLEU-4 score. As shown in Table C.2, our method, SpaMo, achieves the highest BLEU-4 score of 24.32 without the need for the visual fine-tuning. SpaMo requires 22.7M trainable parameters, which is relatively efficient compared to other methods such as GFSLT-VLP (215.6M), Sign2GPT (16M), and Fla-LLM (>705.6M). This demonstrates that SpaMo effectively balances model complexity and training efficiency to achieve superior performance without the additional step of the visual fine-tuning.
C.4 More Qualitative Results
We provide additional translation examples for PHOENIX14T, CSL-Daily, and How2Sign. As shown in Table 14, in PHOENIX14T, our method consistently delivers accurate translations, while GFSLT-VLP struggles to capture the correct semantic meaning.
In CSL-Daily, we present a comparison between glosses and visual tokens, as well as between reference translations and generated translations. As shown in Table D, most visual tokens are matched to the glosses, though they are not perfectly aligned. Notably, in the last three examples, the visual tokens include words that are missing from the glosses but appear in the reference translations.
For How2Sign, Table 16 presents translation results along with their corresponding visual tokens. Since How2Sign lacks gloss annotations, we include actual sign frames for qualitative comparison. Similar to the results on PHOENIX14T and CSL-Daily, many visual tokens in How2Sign are closely aligned with the translations. Note that although OpenSLT Tarrés et al. (2023) is the only publicly available baseline [^6] https://github.com/imatge-upc/slt_how2sign_wicv2023, we were unable to reproduce their results due to a broken link to the fine-tuned I3D features at the time of drafting.
Pairs Cosine Similarity
Glosses & Target Translations 0.8400
Visual Tokens & Target Translations 0.6781
[…]
Table 13: Cosine similarity between the visual tokens and the glosses.
Figure 6: Performance curves across epochs.
C.5 Cosine Similarity Between Visual Tokens and Glosses
We use Sentence-BERT Reimers and Gurevych (2019) to evaluate the similarity between the generated visual tokens and the ground-truth glosses from PHOENIX14T, using cosine similarity. Additionally, we assess the similarity between the visual tokens and the target translation, as well as the alignment between the glosses and the translation, which highlights varying degrees of correspondence.
As shown in Table C.4, and as we expected, the highest similarity occurs between glosses and target translations, indicating a strong semantic correspondence. By contrast, the visual tokens show lower similarity to both target translations and glosses, suggesting that they are not random. The reason for the score not being higher is likely the inclusion of unrelated words in the visual tokens compared to the actual glosses, as illustrated in Table 7.
C.6 Performance Curve Across Epochs
Figure 6 shows the performance curves of SpaMo using BLEU-4 and ROUGE over 40 training epochs on the PHOENIX14T dataset. For comparison, we note that other models such as SignLLM, Sign2GPT, and Fla-LLM are trained for 20, 100, and 75 epochs, respectively. These results highlight the progressive improvements in SpaMo’s performance as training advances, offering a detailed look at its efficiency relative to other models.
Appendix D Feasibility of Existing VLMs in SLT
Recent advancements in Vision-Language Models (VLMs) Alayrac et al. (2022); Li et al. (2023); Liu et al. (2024b); Cheng et al. (2024) have enabled LLMs to comprehend various modalities including images and videos, beyond just text. However, in the SLT domain, current VLM designs are not well-suited for processing long sequences of sign videos. As shown in Table 11, the average sign video length exceeds 116 frames, which is significantly longer than typical action recognition or video-text datasets, where sample lengths are often under 16 frames. For example, Flamingo Alayrac et al. (2022), a widely recognized vision-language model, uses 8 frames during training and only 32 frames during inference—far fewer than what is required for SLT. Similarly, VideoLlama2 Cheng et al. (2024) employs 8 frames for training. Moreover, recent LLM-based SLT methods Wong et al. (2024); Gong et al. (2024), including our method, can be classified within the VLM family but are specialized to process and capture long sign video sequences.
@@ Appendix D Feasibility of Existing VLMs in SLT
{CJK*}
UTF8gbsn
Gloss: 你 小 张 什么 时间 认识
Vis. Token: 三 机场 哪里 什么 时候认识
[…]
Ours: 你 和 小张 什么 时候 认识 的?
Gloss: 椅子 他们 想 什么 时间 去 买
Vis. Token: 椅子 时候 什么 你们 他们 每天 什么 时候 旅游 女儿 考试?
[…]
Ours: 他们 想 什么 时候 去 买 椅子?
Gloss: 不是 他 去 见面 同学
Vis. Token: 不是为什么一个 看见他人看同学桌?
[…]
Ours: 不是, 他 是 去 看 同学 。
Gloss: 这 衣服 红 怎么样 这 是 新
Vis. Token: 孩子 北京衣服红色喜欢我他怎么样认识跑身体跑 个他 北京多?
[…]
Ours: 这 件 红色 的 衣服 怎么样? 是 今年  的 。
Gloss: 中午 吃 好了 要 多 吃 水果
Vis. Token: 饮料 了冷了 了老板安全 粥早 喜欢饮料吃饭 水果 吃饭
[…]
Ours: 吃 完 午饭 我 要 多 吃 点 水果 。
Gloss: 你们 吃 什么 我 请客
Vis. Token: 你们大家来什么他怕十分钟
[…]
Ours: 你们 吃饭 什么? 我 请 你们 吃饭 。
Gloss: 超市 我 要 买 椅子 你 去
Vis. Token: 超市 事 茶 怕 椅子 时候 什么 他 来 去 吗?
[…]
Ours: 我 要 去 超市 买 椅子, 你 去 吗?
Gloss: 我  驾驶  高铁  站  接  儿子
Vis. Token: 大憾怕我我们下雨大学今天
[…]
Ours: 天气 太 冷 了, 我们 去 打 篮球 吧 。
Gloss: 穿  暖  没有  我  想  不  买
Vis. Token: 衣服 饱 热情 急 没有 觉得 想 不是 不是
[…]
Ours: 这 件 衣服 不 保暖, 我 不 想 买 。
Table 15: Translation results on the CSL-Daily test set. Exact visual token matches within glosses are highlighted in green. Words highlighted in blue are not present in the glosses but appear in the translation. Correctly translated 1-grams are shown in blue.